# Krylov fractality and complexity in generic random matrix ensembles

### v1

@@ Abstract
An immediate extension of this method could be the utilization of these probes in systems undergoing localization transitions. The behavior of the density of states (10) shares similar features in two extreme limits of the spin-glass Erdős and Schröder (2014) or the double-scaled Sachdev-Ye-Kitaev model Berkooz et al. (2018). It would be intriguing to assess how the proposed Ansatz (9) compares to the tridiagonal elements found in the transfer matrices Berkooz et al. (2018) of these models with corresponding $q$-Gaussian distribution Erdős and Schröder (2014). Future research might explore such transitions within various matrix ensembles and different diagonal distributions Venturelli et al. (2023). We anticipate that the overarching characteristics will align with our findings. These probes could also be applied to non-Hermitian extensions within the 32 symmetry classes, particularly those lacking a fractal phase in generic complex diagonal Hamiltonians De Tomasi and Khaymovich (2022). In generic systems, extracting the exact form of the Lanczos coefficients across the full spectrum with a known density of states is also unknown. We sketch a general methodology in the SM5 sup that may benefit such endeavors.
@@
Note added: While finishing this manuscript, we became aware of an upcoming work Chakraborty et al. that also obtains the same logarithmic function (8) in the Poisson limit.
@@
Acknowledgements: We thank Tanay Pathak for the initial collaboration and feedback. We also acknowledge valuable discussions with Alexei Andreanov, Adolfo del Campo, Barbara Dietz, Anatoly Dymarsky, András Grabarits, Ivan Khaymovich, Rafael Molina, Horacio Pastawski, Anatoli Polkovnikov, Dario Rosa, Lea Santos, Ruth Shir, and Luca Tessieri. B.B. thanks the organizers of the International Workshop of Disorder Systems (Salamanca) 2024 and the University of Luxembourg for their hospitality during the final stages of the preparation of this manuscript. P.N. thanks the University of Kentucky and the Berkeley Center for Theoretical Physics, University of California, Berkeley, for hosting him through the Adopting Sustainable Partnerships for Innovative Research Ecosystem (ASPIRE) program of the Japan Science and Technology Agency (JST), Grant No. JPMJAP2318, during the final stages of the work. B.B. acknowledges financial support from the Institute for Basic Science (IBS) in the Republic of Korea through Project No. IBS-R024-D1. The work of P.N. is supported by the Japan Society for the Promotion of Science (JSPS) Grants-in-Aid for Transformative Research Areas (A) “Extreme Universe” No. JP21H05190.
@@ Appendix D SM4: Log-variance of the Lanczos spectrum
We further investigate the Krylov spectrum to quantify the two transitions in Krylov space. The probe we consider numerically is the log-variance of the Lanczos spectrum Rabinovici et al. (2022a). This gives us information about the spread of the spectrum around its mean value and is in general understood to capture the ergodic-localized transition. We study its behavior in the multifractal regime and see how that changes from across the two transition points. We consider $12$ system sizes and study how the log-variance, defined as Rabinovici et al. (2022a)
@@
as a function of the parameter $\gamma$. The results presented correspond to averaging over many realizations. It is expected that the variance will be low in the ergodic regime and higher in the non-ergodic regimes Rabinovici et al. (2022a). This is observed for the RP model and is presented in Fig. 7 (left). It is evident that $\sigma_{b}$ nearly vanishes in the ergodic regime $\gamma<1$ and begins to grow upon ergodicity breaking for $\gamma>1$. Naturally, the growth is slower for larger system sizes since the effective ergodicity breaking parameter decreases with increasing system size.
@@
Appendix E SM5: Tridiagonalizing Rosenzweig-Porter model
@@ E.0.3 Householder matrix
E.0.4 Orthogonal transformation revisited
@@
E.0.5 Tridiagonal Rosenzweig-Porter
@@
Appendix F SM6: Krylov-Eigenstate Overlap
@@
The quest is then to solve the three-term recurrence relation for an energy $E_{m}$. For the systems we consider, composed of Wigner matrices, it is known that the mean $\overline{a_{n}}=0$. Likewise it is also known that $\overline{b_{n}}=\sqrt{\frac{2}{\beta N}}\frac{\Gamma((1+(N-n)\beta)/2)}{\Gamma((N-n)\beta/2)}$ Balasubramanian et al. (2023a) in the ergodic phase. We shall use these facts as approximations for the coefficients in (59)

### v2

@@ Abstract
An immediate extension of this method could be the utilization of these probes in systems undergoing localization transitions. The behavior of the DOS (10) shares similar features in two extreme limits of the spin-glass Erdős and Schröder (2014) or the double-scaled SYK model Berkooz et al. (2018) (see SM for some initial attempts). It would be intriguing to assess how the proposed Ansatz (9) compares to the tridiagonal elements found in the models with corresponding $q$-normal distribution Erdős and Schröder (2014); Berkooz et al. (2018). This seems to be illuminating if one computes the moments of the Hamiltonian (1) using the chord diagrams Zhu , where the $A$ and $B$ matrices provide the intersecting and the nonintersecting chords. Future research might explore such transitions within various matrix ensembles and different diagonal distributions Venturelli et al. (2023). We anticipate that the overarching characteristics will align with our findings. These probes could also be applied to non-Hermitian extensions within the 38-fold symmetry classes Kawabata et al. (2019), particularly those lacking a fractal phase in generic complex diagonal Hamiltonians De Tomasi and Khaymovich (2022). In generic systems, extracting the exact form of the Lanczos coefficients across the full spectrum with a known DOS is also unknown. We sketch a general methodology in SM5 that may benefit such endeavors.
@@
Note added: Recently, we became aware of an upcoming work Chakraborty et al. that also obtains the same logarithmic function (8) in the Gaussian limit.
@@
Acknowledgments: We thank Tanay Pathak for the initial collaboration and feedback. We also acknowledge valuable discussions with Alexei Andreanov, Adolfo del Campo, Barbara Dietz, Anatoly Dymarsky, András Grabarits, Ivan Khaymovich, Rafael Molina, Horacio Pastawski, Anatoli Polkovnikov, Dario Rosa, Lea Santos, Ruth Shir, Luca Tessieri, and Zhuo-Yu Xian. B.B. thanks the organizers of the International Workshop of Disorder Systems (Salamanca) 2024 and the University of Luxembourg for their hospitality during the final stages of the preparation of this Letter. P.N. thanks the University of Kentucky and the Berkeley Center for Theoretical Physics, University of California, Berkeley, for hosting him through the Adopting Sustainable Partnerships for Innovative Research Ecosystem (ASPIRE) program of the Japan Science and Technology Agency (JST), Grant No. JPMJAP2318, during the final stages of the work. B.B. acknowledges financial support from the Institute for Basic Science (IBS) in the Republic of Korea through Project No. IBS-R024-D1. The work of P.N. is supported by the Japan Society for the Promotion of Science (JSPS) Grant-in-Aid for Transformative Research Areas (A) “Extreme Universe” No. JP21H05190.
@@ Appendix D SM4: Log-variance of the Lanczos spectrum
We further investigate the Krylov spectrum to quantify the two transitions in Krylov space. The probe we consider numerically is the log-variance of the Lanczos spectrum Rabinovici et al. (2022). This gives us information about the spread of the spectrum around its mean value and is in general understood to capture the ergodic-localized transition. We study its behavior in the multifractal regime and see how that changes across the two transition points. We consider $12$ system sizes and study how the log-variance, defined as Rabinovici et al. (2022)
@@
as a function of the parameter $\gamma$. The results presented correspond to averaging over many realizations. It is expected that the variance will be low in the ergodic regime and higher in the non-ergodic regimes Rabinovici et al. (2022). This is observed for the RP model and is presented in Fig. 7 (left). It is evident that $\sigma_{b}$ nearly vanishes in the ergodic regime $\gamma<1$ and begins to grow upon ergodicity breaking for $\gamma>1$. Naturally, the growth is slower for larger system sizes since the effective ergodicity breaking parameter decreases with increasing system size.
@@
Appendix E SM5: Moments of the distribution (10)
Figure 8: The density of the states (DOS) (26) for different $q$ is shown. For both cases, the extreme limits $q=0$ and $q=1$, match perfectly for the semicircle and the Gaussian distribution.
We briefly discuss the properties of the distribution of the moments of the DOS (10). For convenience, we choose $p=(1-q/2)$, which interpolates between the semicircle and the normal distribution. However, a more general form $p=(1-f(q)/2)$ can be chosen such that $f(0)=0$ and $f(1)=1$. With the above choice, the Lanczos coefficients are given by
$\mathsf{b}(x)^{2}=\left(1-\frac{q}{2}\right)\left(\frac{1-x^{1-q}}{1-q}\right)=-\left(1-\frac{q}{2}\right)\ln_{q}x\,,$ (25)
with $q\equiv q(\gamma)$. For $q=0$ and $q=1$, it reduces to the corresponding semicircle and the Gaussian limits. For the generic form of (25), the DOS is given by
$\rho_{\mathrm{toy}}(E)=\sqrt{\frac{1}{2\pi(2-q)(1-q)}}\frac{\Gamma\big{(}\frac{1}{1-q}\big{)}}{\Gamma\big{(}\frac{1}{2}+\frac{1}{1-q}\big{)}}\Big{(}1-\frac{(1-q)}{2(2-q)}E^{2}\Big{)}^{\frac{1+q}{2(1-q)}}\,.$ (26)
[…]
Due to the symmetry of $\rho_{\mathrm{RP}}(-E)=\rho_{\mathrm{RP}}(E)$, the odd moments vanish. The even moments can be straightforwardly calculated and are given by
$m^{\mathrm{toy}}_{2n}(q)=(2n-1)\text{!!}\,\frac{(2-q)^{n}}{(1-q)^{n+1}}\frac{\Gamma\big{(}\frac{1}{1-q}\big{)}}{\Gamma\big{(}n+1+\frac{1}{1-q}\big{)}}\,.$ (28)
Since $10\leq q<$ 1, the denominator never vanishes. The moments in the two extreme limits are given by
$\lim_{q\rightarrow 0}m^{\mathrm{toy}}_{2n}(q)=\frac{2^{n}(2n-1)\text{!!}}{\Gamma(n+2)}=\frac{(2n)!}{n!(n+1)!}\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \lim_{q\rightarrow 1}m^{\mathrm{toy}}_{2n}(q)=(2n-1)\text{!!}=\frac{(2n)!}{2^{n}n!}\,.$ (29)
From the above moments (28), we employ the iterative algorithm Viswanath and Müller (1994); Parker et al. (2019); Nandy et al. (2024) to compute the lower-order Lanczos coefficients:
$a^{\mathrm{toy}}_{n}(q)=0\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ b^{\mathrm{toy}}_{n}(q)=\sqrt{\frac{(2-q)n\left(\frac{n+1}{2}-\frac{n-1}{2}q\right)}{[(n+1)-nq]\,[n-(n-1)q]}}\,$ (30)
Since $10\leq q<$ 1, the denominator is always non-vanishing. In the two extreme limits, we have
$\lim_{q\rightarrow 0}b^{\mathrm{toy}}_{n}(q)=1\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \lim_{q\rightarrow 1}b^{\mathrm{toy}}_{n}(q)=\sqrt{n}\,.$ (31)
For finite-size systems, such lower-order coefficients can be matched with the numerical Lanczos coefficients with an initial state that has smooth dependence on the energy basis Balasubramanian et al. (2023b). We leave such study for future work.
Appendix F SM6: Tridiagonalizing Rosenzweig-Porter model
@@ F.0.3 Householder matrix
F.0.4 Orthogonal transformation revisited
@@
F.0.5 Tridiagonal Rosenzweig-Porter
@@
It is interesting to note that a consequence of this algorithm is that the mean of the off-diagonal elements remains larger than the standard deviation of the diagonal elements. This can be understood by first noting that $\alpha$ in equations (56)-(59) is $\mathcal{O}(1)$ for the RP model, while $\beta$ is proportional to $1/N^{\gamma}$. The standard deviation of the diagonal elements then scales at most as $1/N$ (without rescaling, near the beginning of the Krylov chain). It is even more strongly suppressed towards the middle of the chain. The mean of the off-diagonal elements (before rescaling) scales as $\sqrt{N}\beta$, leading to a much larger contribution. The $\sqrt{N}$ factor arises because the norm of a Gaussian random vector follows the Nakagami distribution, which has a finite mean. The diagonal elements remain Gaussian distributed.
@@
Appendix G SM7: Krylov-Eigenstate Overlap
@@
The quest is then to solve the three-term recurrence relation for an energy $E_{m}$. For the systems we consider, composed of Wigner matrices, it is known that the mean $\overline{a_{n}}=0$. Likewise it is also known that $\overline{b_{n}}=\sqrt{\frac{2}{\beta N}}\frac{\Gamma((1+(N-n)\beta)/2)}{\Gamma((N-n)\beta/2)}$ Balasubramanian et al. (2023a) in the ergodic phase. We shall use these facts as approximations for the coefficients in (65), along with the underlying assumption that there is no correlation between $b_{n}$ and $\eta^{n}_{m}$.